Skip to main content

IODP Expedition 395E

Complete South Atlantic Transect Reentry Systems

Week 8 Report (23 – 29 May 2021)

Download PDF

Week 8 of Expedition 395E, Complete South Atlantic Transect Reentry Systems, was spent in transit to Reykjavík, Iceland.



At the beginning of Week 8, the ship had completed 2427 nmi of the total 5761 nmi transit to Reykjavík. During the week, the ship transited a further 1850 nmi at an average speed of 11.0 kt. By midnight on 29 May, we had completed 4277 nmi of the transit, with 1484 nmi remaining.

Science Results

Scientific measurements on cores and samples from Sites U1560 and U1561 were already completed before this week.


No onboard Outreach Officer is sailing during Expedition 395P. Social media posts were made via the JR Facebook ( and Twitter ( accounts, run by the JRSO technical staff.

On 28 May, the Expedition 390/393 Co-Chief Scientists and the U.S. Science Support Program Senior Education and Outreach Officer held a Facebook Live event, which was also streamed live on YouTube (

Technical Support and HSE Activities

Laboratory Activities

  • Continued to investigate the P-wave velocity gantry problems:
    • Reset to the initial settings.
    • The rotation of the gear that moves the calipers was set to a lower speed, so that the movement no longer jams.
    • Investigation is ongoing, but the system is operational
  • Organized offgoing shipments.
  • There has been some condensation in the cold room, which is ameliorated when the natural vent and the blanking plate are closed when the cold room is on.
  • Working with Siem Offshore on a port call work plan for limestone delivery and magnetometer cable replacement.
  • Investigation of the precision depth recorder (PDR) 3.5 kHz signal found that it worked at lower power but the signal was lost at higher power.
  • Investigated discrete sample rotations on the superconducting rock magnetometer (SRM). On Expedition 384, results indicated that the code for discrete sample rotations is not correct. Testing is in progress.

IT Support Activities

  • Recovered NEBO (Cumulus) server after a software conflict with a USB device
  • Organized and prepared freight for shipment.
  • Firmware updated on both Secure Mobile Access (SMA) 200 Virtual Private Network (VPN) appliances.
  • Deployed the latest versions of Firefox and Chrome shipwide.

Developer Activities

  • Working on changes to the LORE Thermal Conductivity (TCON) report.
  • Investigating two columns missing from the Gas Elements Standard LORE report.

HSE Activities

  • Safety shower and eye wash stations were tested.